# A novel 3D food printing technique: achieving tunable porosity and fracture properties via liquid rope coiling

### v1

@@ 3.4.7 Sensory anticipation
The key question remaining is whether individuals could distinguish differences between the small fractures that make up a full fracture through the structure. We are not going to address this question in this work; however, we will briefly discuss it based on the available literature studies. Previous research found a general human force perception threshold of 1-6 N, influenced by the direction and speed of mastication and the magnitude of the forces [29]. Peaks below 1 N were filtered out in the current study, so in theory, all reported forces should be perceivable. However, the sensory perception also depends on the product type, where a higher sensitivity has been reported for harder products with significant force differences between fracture peaks, but also depends on the time, kinetics, and sound emissions of the fractures [45, 46, 47, 48]. The ranges of fracture forces we identified offer some insights into this perception but do not address differences between consecutive peaks. Hence, further research should quantify these peak force differences to understand the relationship between a structure’s coiled geometry and its fractural properties. Following such an approach in combination with sensory panel evaluations in future studies will provide a comprehensive understanding of the corresponding sensory properties of coiled structures, including bolus formation and additional sensory cues.
@@ 4 Conclusion
To further expand the capabilities of 3D food printing, we presented a new approach that exploits the liquid rope coiling effect by depositing the filament from a high distance. This method enables the creation of coiled food structures with unique morphological features and tunable textural properties, both primarily driven by the porosity of the coiled structure. The coiled structure was controlled by first identifying optimal printing conditions to induce a specific coiling regime, in this case, translated coiling. The coiling regime could be controlled through the 3D printing parameters, including deposition height, printing speed, and volumetric flow rate. Consequently, this allowed us to control the degree of overlap between coiled filaments and hence the porosity of the food structure. By characterizing the mechanical properties and morphology of the coiled food structures, we demonstrate that the mechanical behavior of coiled structures can be engineered by manipulating their porosity.
In addition, we introduced methods of analyzing fracture data and tried to connect these to the structural properties of the samples. We observed that in a range of porosity, increasing the porosity reduced the hardness and initial stiffness of samples. However, porosity had no major effect on the brittleness, implying that this parameter is mostly dependent on the compositional and microstructural differences among the samples. We observed that the number of fracture events, representing the crunchiness, depends on the composition and consistently is higher for compression tests than for cutting tests. Furthermore, fracture force distributions demonstrated a universality following the Gutenberg–Richter scaling law.
Based on these findings, we have demonstrated that the utilization of liquid rope coiling in 3DFP offers a relatively faster way of food printing and enables tuning textural properties, which reveals the potential of using coiling for further applications of 3DFP in food design and development. To establish connections between our results and the sensory perception further sensory studies are recommended.
@@
Materials
Toasted green pea flour was made by toasting (120°C, with 100% RH super-heated steam for 30 minutes) wrinkled green peas (Pisum sativum L.) in a Rational convection oven (SSC WE 61). Subsequently, the toasted peas were dehulled and ground in a ZPS 100 mill (Hosokawa-Alpine) at a milling speed of 20000 rpm, after a speed of 4000 rpm, and air throughput of 50 m3/h, which obtained a flour with a D50 of 24 $\mu$m and D90 of 80 $\mu$m. Carrot powder consisted of dried and milled carrot pulp, as a residue after carrot juice extraction. This food ingredient was kindly supplied by Van Rijsingen Green, Helmond (The Netherlands).
@@
In this formula, $h_{\text{layer}}$ is the layer height equal to $d/2$, $w_{\text{track}}$ is the track width equal to $3\cdot h_{\text{layer}}$, and $A_{\text{syringe}}$ is the circular area of the syringe body. The volumetric flow rate ($Q$) describes the volume of material deposited from the nozzle per unit of time. It is dependent on the speed of injection and the area from which it is ejected and can be expressed as:
@@
Furthermore, the flow rate is also directly correlated to the macrostructural porosity of the coiled sample, as can be observed in Supplementary Figure 1. Here it can be observed that higher flow of the ink from the nozzle will result in 3DFP structures with lower porosities, and vice versa for low flow rates.
@@
Rheology of the inks
Ink printability and flow behavior
Besides $\sigma_{\text{yield}}$ (Supplementary Table 1), the flow behavior index ($n$) and consistency index ($K$) can also be retrieved from the Herschel-Bulkley model, and are provided in Supplementary Table 2, along with $\sigma_{\text{yield}}$. Although these indices are not commonly reported in the literature, low $n$ and $K$ values are associated with shear-thinning properties, whereas higher values indicate that materials cannot be extruded easily [31, 49]. For all inks, $n<1$ holds true, signifying a shear-thinning behavior and good extrudability.
@@
In order to validate that we dealt with liquid rope coiling (LRC) rather than solid rope coiling (SRC) in this study, the effective bending stress ($B_{\text{eff}}$) and effective viscosity of the inks were established by using the formulas presented below for prediction of the coiling frequency of liquid and solid materials in the gravitational regime. This $B_{\text{eff}}$ gives an indication of the stress in the filament during extrusion.
@@
Ink stability
In addition to being able to pass through a nozzle and to exhibit coiling behavior, the ink properties should also allow for a fast stabilization once the ink has been deposited, to ensure that the printed coiled structure does not collapse [49]. This post-printing stability is mostly dependent on the ink’s viscoelastic moduli, $G^{\prime}$ and $G^{\prime\prime}$, and loss tangent $\tan\delta$ under large oscillatory shear (LOAS) ranges, as these indicate how much stress a material can withstand prior to collapsing. From Figure 2d (main text), it can be observed that $G^{\prime}$ remains nearly constant in the linear viscoelastic regime (LVR), regardless of ink type, until a critical strain is reached, after which it falls abruptly. This indicates the transition to the nonlinear viscoelastic regime (NLVR), where the material’s network starts to break down, resulting in shear-thinning behavior [50]. In the LVR, inks have $G^{\prime}$ values greater than the minimum of 300 Pa reported in literature for post-printing stability [31], so therefore are guaranteed to yield stable prints. Moreover, the $\tan\delta$ of the inks in the LVR are 0.18, 0.13, and 0.36 for pea, carrot, and cookie, respectively. The former two fit within a range of $0.052\leq\tan\delta\leq 0.268$, which was found to deliver stable prints [51, 52, 53]. The $\tan\delta$ of the cookie was closer to one, indicating that it is closer to exhibiting flow behavior, and would, therefore, be less self-supporting. However, the previously indicated range was established for food inks with very different compositions, so may not act as an accurate representation for the inks used in this study.
@@
Evaluation of Lissajous plots and nonlinear response upon LAOS
Additional Lissajous plot analyses and quantifications from the LAOS deformation are able to provide more details on the inks during their transition from the LVR to the NLVR region, which is also what occurs during extrusion from the nozzle during 3DFP. This type of more in-depth rheological analysis through the use of Lissajous plots of food-based inks have only recently started to gain attention by scientists [54, 55, 56, 57]. From the elastic Lissajous plots in Supplementary Figure 3a, it can be observed that at 0.01% strain, all inks exhibit a linear viscoelastic response, characterized by their narrow elliptical shape. As the imposed strain amplitude increases from 1 to 10%, the ellipses start to widen as a result of $G^{\prime}$ decreasing faster than $G^{\prime\prime}$ (Supplementary Figure 3b). Ultimately, the data shifts to a parallelogram shape, implying a shift to a nonlinear viscoelastic response; i.e., from a solid-dominated to a liquid-dominated behavior [58, 50]. This is also what occurs during the extrusion of ink from the nozzle. Furthermore, the largest change in shape under large strain amplitudes is observed for cookie and pea inks, which indicates a greater structural breakdown of the inks under large deformation, although complete microscopic fracture did not occur within the strain range tested.
Supplementary Figure 3: (a) Top three rows: the elastic Lissajous plots for pea (green), carrot (orange), and cookie (yellow) inks. (b) Bottom row-left: Viscoelastic moduli versus shear strain; $G^{\prime}$=solid symbols, $G^{\prime\prime}$=open symbols. (c) Bottom row-middle: Stiffening ratios. (d) Bottom row-right: Dissipation ratio; $\sigma$=stress, $\gamma$=strain.
@@
Further information on coiling
Changing the flow rate through modification of EM and Uxy
@@
The effects of printing parameters on coiling behavior
@@
Printing parameters used for the creation of the coiled structures
See Supplementary Figure 8.
@@
Porosity of the coiled structures
@@
Texture analysis
Fracture behavior of bulk materials
@@
Initial stiffness of coiled pea structures
The initial stiffness of samples upon cutting and compression is shown in Supplementary Figure 10.

### v2

@@ 3.4.7 Sensory anticipation
The key question remaining is how individuals would perceive the sensory response of the coiled structures and would distinguish between various fracture profiles. Addressing this question remains beyond the scope of this research; however, we will briefly discuss it based on the available literature studies. Previous research found a general human force perception threshold of 1-6 N, influenced by the direction and speed of mastication and the magnitude of the forces [33]. Peaks below 1 N were filtered out in the current study, so in theory, all reported forces should be perceivable. However, the sensory perception also depends on the product type, where a higher sensitivity has been reported for harder products with significant force differences between fracture peaks, but also depends on the time, kinetics, and sound emissions of the fractures [47, 48, 49, 50]. The ranges of fracture force we identified offer some insights into this perception but do not address differences between consecutive peaks. Hence, further research should quantify these peak force differences to understand the relationship between a structure’s coiled geometry and its fracture properties. Following such an approach in combination with sensory panel evaluations in future studies will provide a comprehensive understanding of the corresponding sensory properties of coiled structures, including bolus formation and additional sensory cues.
@@ 4 Conclusion
To further expand the capabilities of 3D food printing, we presented a new approach that exploits the liquid rope coiling effect by depositing the filament from a high distance. This method enables the creation of coiled food structures with unique morphological features and tunable textural properties, both primarily driven by the porosity of the coiled structure. The coiled structure was controlled by first identifying optimal printing conditions to induce a specific coiling regime, in this case, translated coiling. The coiling regime could be controlled through the 3D printing parameters, including deposition height, printing speed, and volumetric flow rate. Consequently, this allowed us to control the degree of overlap between coiled filaments and hence the porosity of the food structure. By characterizing the mechanical properties and morphology of the coiled food structures, we demonstrated that the mechanical behavior of coiled structures can be engineered by manipulating their porosity based on the printing parameters.
In addition, we introduced methods of analyzing fracture data and tried to connect these to the structural properties of the samples. We observed that in a range of porosities, increasing the porosity reduced the hardness and initial stiffness of samples. However, porosity had no major effect on the brittleness, implying that this parameter is mostly dependent on the compositional and microstructural differences among the samples. We observed that the number of fracture events, representing the crunchiness, depends on the composition and consistently is higher for compression tests than for cutting tests. Furthermore, fracture force distributions demonstrated a universality following the Gutenberg–Richter scaling law.
Based on these findings, we have demonstrated that the utilization of liquid rope coiling in 3DFP offers a relatively faster way of food printing and enables tuning textural properties, which reveals the potential of using coiling for further applications of 3DFP in food design and development.
Acknowledgements
We would like to express our gratitude to Mokhtar Adda-Bedia for the insightful discussions. We also thank Gerard Gimènez-Ribes for his support in analyzing and interpreting the rheology data.
Author Contributions Statement
A.G. and S.J.G. Contributed equally to this work. A.G. M.N. and M.H. supervised and designed the research. S.J.G. conducted the experiments and wrote the original draft. S.A.G. helped with and supervised the 3D printing-related procedures. A.G. and S.J.G performed the data analysis, visualizations, and revising the manuscript. A.G. developed data analysis codes. M.H. acquired funding for the project. M.H., M.N., and S.A.G. contributed to the validation. All authors discussed the results and contributed to establishing the methodology and experimental procedures. All authors have reviewed the results and approved the final version of the manuscript.
@@ Acknowledgements
SI Detailed information on Materials and Methods
SI.1 Materials
Toasted green pea flour was made by toasting (120°C, with 100% RH super-heated steam for 30 minutes) wrinkled green peas (Pisum sativum L.) in a Rational convection oven (SSC WE 61). Subsequently, the toasted peas were dehulled and ground in a ZPS 100 mill (Hosokawa-Alpine) at a milling speed of 20000 rpm, zifter speed of 4000 rpm, and air throughput of 50 m3/h, which obtained a flour with a D50 of 24 $\mu$m and D90 of 80 $\mu$m. Carrot powder consisted of dried and milled carrot pulp, as a residue after carrot juice extraction. This food ingredient was kindly supplied by Van Rijsingen Green, Helmond (The Netherlands).
@@
In this formula, $h_{\text{layer}}$ is the layer height equal to $d/2$, $w_{\text{track}}$ is the track width equal to $3\cdot h_{\text{layer}}$, and $A_{\text{syringe}}$ is the circular area of the syringe body. The volumetric flow rate ($Q$) describes the volume of material deposited from the nozzle per unit time and is dependent on the speed of injection and the area from which it is ejected and can be expressed as:
@@
Furthermore, the flow rate is also directly correlated to the macrostructural porosity of the coiled sample, as can be observed in 1. Here it can be observed that higher flow of the ink from the nozzle will result in 3DFP structures with lower porosities, and vice versa for low flow rates.
@@
SI Further information on the rheological properties of the 3DFP inks
SI.1 Ink printability and flow behavior
Besides $\sigma_{\text{yield}}$ (Supplementary Table 1), the flow behavior index ($n$) and consistency index ($K$) can also be retrieved from the Herschel-Bulkley model, and are provided in Supplementary Table 2, along with $\sigma_{\text{yield}}$. Although these indices are not commonly reported in literature, low $n$ and $K$ values are associated with shear-thinning properties, whereas higher values indicate that materials cannot be extruded easily [35, 51]. For all inks, $n<1$ holds true, signifying a shear-thinning behavior and good extrudability.
@@
In order to validate that we dealt with liquid rope coiling (LRC) rather than solid rope coiling (SRC) in this study, the effective bending stress ($B_{\text{eff}}$) and effective viscosity of the inks were established by using the formulas presented below for prediction of the coiling frequency of liquid and solid materials in the gravitational regime [36, 12, 37]. This $B_{\text{eff}}$ gives an indication of the stress in the filament during extrusion.
@@
SI.2 Ink stability
In addition to being able to pass through a nozzle and to exhibit coiling behavior, the ink properties should also allow for a fast stabilization once the ink has been deposited, to ensure that the printed coiled structure does not collapse [51]. This post-printing stability is mostly dependent on the ink’s viscoelastic moduli, $G^{\prime}$ and $G^{\prime\prime}$, and loss tangent $\tan\delta$ under large oscillatory shear (LOAS) ranges, as these indicate how much stress a material can withstand prior to collapsing. From Figure 2d (main text), it can be observed that $G^{\prime}$ remains nearly constant in the linear viscoelastic regime (LVR), regardless of ink type, until a critical strain is reached, after which it falls abruptly. This indicates the transition to the nonlinear viscoelastic regime (NLVR), where the material’s network starts to break down, resulting in shear-thinning behavior [52]. In the LVR, inks have $G^{\prime}$ values greater than the minimum of 300 Pa reported in literature for post-printing stability [35], so therefore are guaranteed to yield stable prints. Moreover, the $\tan\delta$ of the inks in the LVR are 0.18, 0.13 and 0.36 for pea, carrot and cookie, respectively. The former two fit within a range of $0.052\leq\tan\delta\leq 0.268$, which was found to deliver stable prints [53, 54, 55]. The $\tan\delta$ of cookie was closer to one, indicating that it is closer to exhibiting flow behavior, and would, therefore, be less self-supporting. However, the previously indicated range was established for food inks with very different compositions, so may not act as an accurate representation for the inks used in this study.
@@
SI.3 Evaluation of Lissajous plots and nonlinear response upon LAOS
Additional Lissajous plot analyses and quantifications from the LAOS deformation are able to provide more details on the inks during their transition from the LVR to the NLVR region, which is also what occurs during extrusion from the nozzle during 3DFP. This type of more in-depth rheological analyses through the use of Lissajous plots of food-based inks have only recently started to gain attention by scientists [56, 57, 58, 59]. From the elastic Lissajous plots in Supplementary Figure 3a, it can be observed that at 0.01% strain, all inks exhibit a linear viscoelastic response, characterized by their narrow elliptical shape. As the imposed strain amplitude increases from 1 to 10%, the ellipses start to widen as a result of $G^{\prime}$ decreasing faster than $G^{\prime\prime}$ (Supplementary Figure 3b). Ultimately, the data shifts to a parallelogram shape, implying a shift to a nonlinear viscoelastic response; i.e., from a solid-dominated to a liquid-dominated behavior [60, 52]. This is also what occurs during the extrusion of ink from the nozzle. Furthermore, the largest change in shape under large strain amplitudes is observed for cookie and pea inks, which indicates a greater structural breakdown of the inks under large deformation, although complete microscopic fracture did not occur within the strain range tested.
Supplementary Figure 3: (a) The elastic Lissajous plots for pea (green), carrot (orange), and cookie (yellow) inks. (b) Viscoelastic moduli versus shear strain; $G^{\prime}$=solid symbols, $G^{\prime\prime}$=open symbols. (c) Stiffening ratios. (d) Dissipation ratio; $\sigma$=stress, $\gamma$=strain.
@@
SI Further information on coiling
SI.1 Changing the flow rate through modification of EM and Uxy
@@
SI.2 The effects of printing parameters on coiling behavior
@@
SI.3 Printing parameters used for the creation of the coiled structures
See Supplementary Figure 8
@@
SI.4 Porosity of the coiled structures
@@
SI Texture analysis
SI.1 Fracture behavior of bulk materials
@@
SI.2 Fracture of a single filament
Cutting a single filament allows us to quantify the minimum meaningful fracture force. The highest fracture force response of the filament is equivalent to the minimum meaningful fracture in our coiled structures. Supplementary Figure 10, shows the applied force as a function of the applied strain for cookie filaments.
Supplementary Figure 10: Cutting force response of cookie filaments with diameters of (a) 1 mm and (b) 1.5 mm, shown in three repetitions.
SI.3 Initial stiffness of coiled pea structures
The initial stiffness for different porosities is presented in Supplementary Figure 11. The initial stiffness shows considerable variation among the samples; however, its order of magnitude consistently falls within the range of $10^{6}\text{Pa}$ during compression and cutting tests.